OpenStudy (kulm-5):

2x + 4y = 12 y = x – 3 What is the solution to the system of equations? (–1, 8) (8, –1)

jhonyy9 (jhonyy9):

2x+4y=12 divide both sides by 2 and will get x+2y=6 y=x-3 => x-y = 3 x+2y=6 subtract second from first equation x-y=3 ------------ 0 3y = 3 y= 3/3 = 1 y = 1 1=x-3 x = 4 y=1 but this is interesting bc. (4,1) not is there between this choices

jhonyy9 (jhonyy9):

and than you check these given choices so these not are true ,right than you substitute these value of x and y in place of x and y inside equations

jhonyy9 (jhonyy9):

but substituted 4 and 1 will get it correct sure please try,check these